If a^2+b^2+c^2=2a  and  a+b+c=0 find ab+bc+ac 
 
Similar questions solutions

Grade:9

1 Answers

Arun
5 years ago
Dear studnt
 
(a + b + c)^2 = a^2 + b^2 + c^2 + 2 (ab + bc + ca)
 
0 = a^2 + 2 (ab + bc + ca)
 
Hence
 
ab + bc + ca = – a^2 / 2
 
Regards
Arun